OpenStudy (anonymous):

three cars start simultaneously from ajmer to banaras along the same highway.the second car travelled with a speed 10km/hr higher than the first car's speed and arrived at benaras 1 hour earlier than the first car.the third car arrived at benaras 33.33 minutes earlier than the first car,travelling half the time at the speed of first car and other half at the speed of second car.find the total distance covered by these three cars during the journey

OpenStudy (anonymous):

For these questions the hard part is to set the equations. This question requires basic understanding of mechanics too. d=v*t (d is distance, v is velocity/speed and t is time.) To set the equations you have to decide what x,y,z should be. (maybe you will even need more variables, maybe less) Can you give that a try?

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

i did try making equations but unable to reach the answer. the number of unknowns exceed the no. of equations

OpenStudy (anonymous):

OK lets do this :) We will definitely have one unknown, what the question asks us, the distance d. We should call x the speed of the first car. Lets call t the time taken for the first car to arrive from A to B. Lets see what we can do with these so far. Equation 1, d=xt Equation 2, (this is for the second car) d=(x+10)(t-1) This means that it traveled the same distance d, with +10km/h speed and it took -1h Equation 3, (for the 3rd car) Now this one is the trickiest as the question states that this car has changed its speed during the travel. When you read the question an alarm bell should ring in your head! Everything so far was given in hours and now we have 33.33 minutes !!!!!! Thus we have to change it to hours. 33.33min=0.5555 hour We know that it took this car t-0.5555 to do the full journey. For (t-0.5555)/2, that is half the time of this cars travel, it had the same speed as the first car =x. For the other half the time (t-0.5555)/2 with the speed of the second car =x+10 So finally we arrive to our 3rd equation d=(t-.05555)/2 *x + (t-0.5555)/2 * (x+10)

OpenStudy (anonymous):

That is 3 equations with 3 unknowns, the rest is just algebra.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I did not get 600. Do the algebra and show your work to your teacher. I cannot see any mistakes made. Maybe there is but understanding a problem and modeling a situation is far more important than the actual data and the result. Sorry I am not gonna put more effort into this problem. Already spent more than intended

OpenStudy (anonymous):

33.33min this is a bit fishy to be honest. There might be a typo error there